OpenStudy (anonymous):

Evaluate the series 1 + 0.1 + 0.01 + . . .

OpenStudy (anonymous):

It is in geometric progression...so,sum to infinite terms in GP is a/(1-r) where a= first term r = second term / first term = 0.1 /1 =0.1 therefore,sum = a / (1-r) = 1/ (1-0.1) = 1/0.9 = 10 / 9

OpenStudy (anonymous):

So, the answer is 10/9?

OpenStudy (anonymous):

1.11?

OpenStudy (anonymous):

yes...it's actually 1.1111111111...so on..that's 1.1 recurring

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

ok